What is the magnitude of force required to pull a block of massm= 100kgup the incline of 32 degrees with constant speed?
1
Expert's answer
2020-10-25T18:26:34-0400

"F_{gravitational}= mass\times gravity."


"= 100\times 9.8= 980N"


"F_{parrallel}=mg\sin \theta= 980\sin(30\degree)= 460N"


"F_{perpendicular}=mg\cos \theta= 980 \cos(30\degree)= 852.6N"


The forces directed perpendicular to the incline balance each other. Thus Fnorm is equal to Fperpendicular.


"F_{Norm}= 852.6N"


There is no other force parallel to the incline to counteract the parallel component of gravity. Thus, the net force is equal to the Fparallel value.


Thus "F_{Net}= 460N"